OpenStudy (anonymous):

The perimeter of a rectangle is 25 cm, and its dimensions are tripled. What is the perimeter of the larger rectangle? 25 cm 60 cm 75 cm 50 cm

OpenStudy (anonymous):

Rectangle has two dimensions Length and Width.. What you will get when they both are tripled??

OpenStudy (anonymous):

6

OpenStudy (anonymous):

See if L is the length and W is the width then on tripling what you will get???

OpenStudy (anonymous):

60

OpenStudy (anonymous):

Tripling means multiplication by 3.. What do you get when you multiply L with 3?? And what will you get when you multiply W with 3??

OpenStudy (anonymous):

75

OpenStudy (anonymous):

3*L = ??? Perimeter given is 25.. L and W are unknown... 3*L = ??? 3*W = ??